OpenStudy (anonymous):

if the square of a number increased by the number is 42, then what is the number?

OpenStudy (across):

\[x^2+x=42,\]\[x^2+x-42=0.\]

OpenStudy (across):

You could use the quadratic formula here.

myininaya (myininaya):

or factor! :)

myininaya (myininaya):

-42=-6(7) 1=-6+7

OpenStudy (anonymous):

solving with quadratic formula, I get -7, 6.